Show that the angle between the two lines in which the plane x−y+2z = 0 intersects the cone

x2 +y2 −4z2 +6yz = 0 is tan −1 (√ 6 /7)


1
Expert's answer
2021-06-08T09:01:10-0400

Answer:-

The lines of intersection have equation

"\begin{cases}
x^2+y^2-4z^2+6yz=0; \\
 x-y+2z=0; \\
z=t,
\end{cases}"

where 𝑡 ∈ ℝ.

The system  "\begin{cases}
x^2+y^2-4z^2+6yz=0; \\
 x-y+2z=0; \\
z=t,
\end{cases}"

is equivalent to "\begin{cases}
x^2+y^2-4t^2+6yt=0; \\
 x=y-2t; \\
z=t,
\end{cases}"


or "\begin{cases}
y^3-4yt+4t^2+y^2-4t^2+6yt=0; \\
 x=y-2t; \\
z=t,
\end{cases}"


or "\begin{cases}
y(y+t)=0; \\
 x=y-2t; \\
z=t,
\end{cases}"

we obtain two line

"\begin{cases}
x=-2t; \\
 y=0; \\
z=t,
\end{cases}" and "\begin{cases}
x=-3t; \\
 y=-t; \\
z=t,
\end{cases}"

where 𝑡 is real parameter. The direction vector of the first line is 𝑣1 = (−2,0,1) and the direction vector of the second line is 𝑣2 = (−3, −1,1). The angle between the two lines is equal to the angle between the direction vectors of this two lines.


cos(𝑣̂1, 𝑣2 ) ="\frac{v_1.v_2}{|v_1||v_2|}=\frac{6+1}{\sqrt{5}\sqrt{11}}=\sqrt{\frac{49}{55}}"

tan2(𝑣̂1, 𝑣2 ) ="\frac{1}{cos^2(v_1.v_2)}-1=\frac{55}{49}-1=\frac{6}{49}"

The (𝑣̂1, 𝑣2 ) = arctan"√6 \over7" , as cos(𝑣̂1, 𝑣2 ) is positive. Therefore, the angle between two lines is equal to arctan "√6\over 7" =  tan −1 ("√ 6 \over7" )